in one day, a store sold 8/5 as many dvds as cds. The total number of dvds and cds sold that day was 273. How many dvds where sold?
Answer by user_dude2008(1862)

c = cds
d = dvds

d=(8/5)c
c+d=873

c+(8/5)c=273
(13/5)c=273
13c=1365
c=1365/13
c=105

d=(8/5)c=(8/5)(105)=8*21=168

Answers: c=105, d=168